OpenStudy (anonymous):

an alpha particle moving at a speed of 5.20x10^4 m/s enters a perpendiclar magnetic field with a strength of 3.0 mT. The magnitude of the acceleration of the particle as it enters the magnetic field will be a.bc x 10^d m/s^2 ?

OpenStudy (anonymous):

 F magnetic=q(v(vector)cross B(vector))=2*1.6*10^-19(5.2*10^4*3*10^-3) =49.8*10^-18N hence acceleration=49.8*10^-18/4*1.66*10^-27=7.5*10^9 m/sec^2 i.e a=7 b=5 c=0 d=9